Shift report 20011122 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Nov 22 2001 - 16:01:08 EST

  • Next message: Stephen J. Sanders: "draft format update"

    ------------------------------------------------------------
    Date: Thu Nov 22 16:01:08 2001, Shift: 08:00-16:00
    Supervisor: rw, with: 
    Log book#11 pages 23 - 24
    
    Shift Summary:
    ------------------------------------------------------------
    1. Two beam storage was made during the shift. The first one
    was cogged at 8:15. Blue/Yellow=27/33 ZDC=600. The beam was 
    dumped at 13:15 after one hour gap cleaning with low rate.
    The second beam was cogged at 14:25 with the ZDC rate of 2700.
    5 min. later the rate went down to 2200 and 10min. later to 1700.
    T3 currents at the beginning were 75, 65, 60 micro amps. 20min.
    later Yellow ring lost 30% of its beam suddenly and the rate 
    went down to 900. After that the beam became more or less stable.
     
    2.	run 	trigger 6     	Setting
    	5939	170K		FS 8deg-2/3A  MRS same sa before
    	5940	 32K		FS 8deg-1/8A but wrong D3,D4
    				current setup.
    	5941     15K            same as above
    
    	5942	TPC Pedestal run
    	5943	DVM
    	5944	pedestal run
    
    	5945	106K		FS 8deg-1/8A with right current
    				for D3,D4. MRS same
    	5946    100K		same as above
    	5947	running		same as above
    
    3. No new problems have been detected.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Nov 22 2001 - 16:01:56 EST